1. How many moles of O2 are originally in a balloon if the initial volume is 1.00 L and the final volume is 0.67 L.  The final number of moles in the balloon is 8.0 moles.
1
Expert's answer
2021-02-10T01:32:28-0500

According to the ideal gas law:

pV = nRT

where p- pressure, V - volume, n - number of moles, R - gas constant, T - temperature.

As temperature and pressure are constant:

V/n = RTp

From here:

V1/n1 = V2/n2

Finally:

n1 = V1n2 / V2 = 1.00 L × 8.0 mol / 0.67 L = 12 mol


Answer: 12 mol